# Surveying the MLLM Landscape: A Meta-Review of Current Surveys

### v1

@@ 4.3.1 Core Evaluation Domains
This domain evaluates how well MLLMs interpret multimodal inputs, such as text, images, and audio, and integrate information across modalities. Benchmarks in this category include tasks like object detection, scene understanding, and feature extraction. For example, VQA datasets like VQAv2 are foundational for evaluating these abilities but are limited by biases that can inflate model performance [55].
@@
Higher-level capabilities, such as logical reasoning, problem-solving, and multimodal reasoning, are captured in this domain. Tasks that require more sophisticated reasoning across modalities, like visual question answering with complex scenarios, are included in this category. This domain tests the models’ deeper understanding of the relationships between the modalities [56].
@@ 4.3.2 Advanced Evaluation Areas
Models need to be evaluated for their robustness, particularly when faced with adversarial prompts or out-of-distribution data. The benchmarks in this category assess how well MLLMs perform under conditions that simulate real-world challenges. Robustness is also critical in ensuring the models’ safety, especially when deployed in domains like autonomous driving or healthcare [57, 58]. Furthermore, this area assesses how models manage hallucinations and mitigate biases.
@@
This category includes evaluations focused on specialized domains, such as medical image interpretation or legal text analysis, where models must combine general multimodal abilities with domain-specific knowledge. Benchmarks like TallyQA, which test complex reasoning in specific domains, fall into this category [56].
@@ 4.3.3 Task-Specific Benchmarks
Early datasets, such as VQAv2, have been widely used to evaluate MLLMs but exhibit limitations due to their susceptibility to language bias and lack of task complexity. In contrast, more recent datasets like TDIUC and DVQA are designed to evaluate models on fine-grained visual understanding, reasoning, and OCR tasks. These datasets provide a more rigorous assessment of the models’ capabilities [55].
@@
This category focuses on evaluating the interaction between modalities, which is key in tasks like multimodal dialogue or visual question answering that require reasoning across both text and images. Advanced datasets, such as VQDv1, which require complex reasoning and the identification of multiple objects in visual contexts, push models to demonstrate a deeper understanding of multimodal relationships [55].
@@ 4.3.4 Ethical and Societal Implications
This domain addresses the importance of ensuring that MLLMs do not perpetuate or amplify societal biases. Models must be evaluated on their ability to perform equitably across demographic groups, and benchmarks in this category assess the fairness of the models [59].
@@
Ensuring trust in MLLMs is critical, especially in applications where safety is paramount. This category evaluates whether models produce harmful or misleading content and assesses their reliability in sensitive domains. It also evaluates how MLLMs handle uncertainty and avoid hallucinations [60, 61].
@@ 4.4 Efficiency and Adaptation
The surveys on resource-efficient Multimodal Large Language Models (MLLMs) explore various approaches aimed at reducing computational costs while maintaining performance. Xu et al. [62] and Jin et al. [63] emphasize the growing need for MLLMs to be more accessible, particularly in resource-constrained environments like edge computing. Both surveys provide comprehensive taxonomies, covering advancements in architectures, vision-language integration, training methods, and benchmarks that optimize MLLM efficiency. Key methods discussed include vision token compression, parameter-efficient fine-tuning, and the exploration of transformer-alternative models like Mixture of Experts (MoE) and state space models. These efforts collectively aim to balance computational efficiency with task performance, driving MLLM adoption across various practical applications.
On the other hand, Liu et al. [64] address the challenge of adapting MLLMs to specific tasks with limited labeled data. The survey categorizes approaches into prompt-based, adapter-based, and external knowledge-based methods, which help these models generalize better in fine-grained domains such as medical imaging and remote sensing. Few-shot adaptation techniques, such as visual prompt tuning and adapter fine-tuning, are critical for extending the usability of large multimodal models without relying on extensive labeled datasets. Despite advancements, these surveys highlight ongoing challenges, including domain adaptation, model selection, and the integration of external knowledge. Both fields point toward a future where MLLMs are not only more efficient but also more flexible and adaptive in handling diverse, real-world tasks.
@@ 4.5 Data-centric
Bai et al. [65] presents a comprehensive survey on the role of data in the development of Multimodal Large Language Models (MLLMs). The paper emphasizes the importance of the quality, diversity, and volume of multimodal data, which includes text, images, audio, and video, in training MLLMs effectively. It identifies significant challenges in multimodal data collection, such as data sparsity and noise, and explores potential solutions like synthetic data generation and active learning to mitigate these issues. The authors advocate for a more data-centric approach, where the refinement and curation of data take precedence, ultimately improving model performance and advancing MLLM capabilities in an increasingly complex landscape.
Yang et al. [66] investigates the intersection of large language models (LLMs) and code, framing code as a vital tool that enhances LLMs’ capacity to operate as intelligent agents. The paper discusses how code empowers LLMs in tasks such as automation, code generation, and software development. It also tackles key challenges around code correctness, efficiency, and security, which are essential when deploying LLMs in practical applications. By analyzing the integration of code with LLMs, the authors showcase the potential of these models to evolve into autonomous agents capable of managing complex tasks across various domains, thereby broadening the scope and impact of LLMs in AI-driven innovations.
@@ 4.6 Contunual Learning
Feng et al. [67] provides a thorough overview of how Large Language Models (LLMs) are being integrated with external knowledge to enhance their capabilities. The survey categorizes the integration methods into two main approaches: knowledge editing and retrieval augmentation. Knowledge editing involves modifying the input or the model itself to update the outdated or incorrect information, while retrieval augmentation fetches external information during inference without altering the model’s core parameters. The authors present a taxonomy covering these methods, benchmarks for evaluation, and applications such as LangChain and ChatDoctor, which leverage these strategies to address domain-specific challenges. Additionally, the paper explores the handling of knowledge conflicts and suggests future research directions for improving LLM performance in complex, real-world tasks through better integration of multi-source knowledge.
Shi et al. [68] offer an extensive survey on the continual learning (CL) of Large Language Models (LLMs), addressing the critical challenges and methodologies in this field. presents an extensive overview of the challenges and techniques related to the continual learning (CL) of Large Language Models (LLMs). The authors focus on two primary directions of continuity: vertical continual learning (adapting models from general to specific domains) and horizontal continual learning (adapting models over time across various domains). They discuss the problems of "catastrophic forgetting," where models lose knowledge of previous tasks, and the complexity of continually updating models to maintain performance on both old and new tasks. The survey outlines key CL methods, including continual pre-training, domain-adaptive pre-training, and continual fine-tuning. It also evaluates various CL techniques, such as replay-based, regularization-based, and architecture-based methods, to mitigate forgetting and ensure knowledge retention. The authors call for more research into evaluation benchmarks and methodologies to counter forgetting and support knowledge transfer in LLMs, making this an underexplored yet crucial area of machine learning research
@@ 4.7 Evaluation Benchmarks
Lu et al. [55] tackle the key challenges in evaluating multimodal large language models (MLLMs), focusing on the unique complexities these models present. They propose a taxonomy of evaluation metrics designed specifically for multimodal tasks, such as cross-modal retrieval, caption generation, and visual question answering (VQA). The authors highlight how current evaluation frameworks often fail to capture the nuances of multimodal interactions, suggesting the need for more tailored metrics to address these limitations.
Li and Lu [69] provide a comprehensive overview of benchmarking datasets and performance metrics in MLLMs. They argue that the absence of standardized protocols across different modalities hinders the fair comparison of models. Their work calls for the establishment of consistent evaluation frameworks that ensure reliable and equitable performance assessment across varied multimodal tasks.
Chang et al. [57] examine the evaluation methodologies for large language models (LLMs), emphasizing the importance of moving beyond task-specific performance metrics like knowledge reasoning. They point out that many existing benchmarks, such as HELM and BIG-Bench, overlook critical issues such as hallucinations, fairness, and societal implications. The authors advocate for more holistic evaluation practices that take into account not only the technical capabilities of LLMs but also their trustworthiness, robustness, and ethical alignment in real-world applications. Their work underscores the need to assess both the technical and societal impacts of these models to ensure responsible AI deployment.
@@ 4.8 Agents and Autonomous Systems
Xi et al. [70], Wang et al. [71] and a few more papers[72, 73, 62] provide comprehensive surveys on the development of LLM-based autonomous agents, highlighting the key modules and frameworks that form the foundation of these systems. At their core, autonomous agents leveraging large language models (LLMs) rely on four essential components: perception, memory, planning, and action. These modules work in synergy to enable agents to perceive their environment, recall previous interactions, and plan and execute actions in real-time. As Xi et al. [70] describe, this architecture allows agents to be highly adaptable across a variety of domains, from digital assistants to autonomous vehicles. Wang et al. [71] further emphasize the importance of expanding the perception-action loop by incorporating multimodal inputs, ensuring agents can effectively handle complex real-world scenarios, such as those found in industrial automation and gaming.
@@
Despite these advancements, LLM-based agents face several critical challenges. Both surveys identify knowledge boundaries as a significant issue, where agents are constrained when operating in specialized or underexplored domains. Prompt robustness is another challenge, as even minor changes in prompts can lead to unpredictable or erroneous behavior, including hallucinations, where agents generate false information. Shi et al. [68] further explore the challenge of catastrophic forgetting, where agents fail to retain knowledge from previous tasks after being updated with new information. Addressing these challenges requires ongoing improvements in how agents interact with external tools, refine prompt structures, and manage memory to prevent knowledge decay and ensure consistent behavior.
@@
LLM-based agents have demonstrated significant versatility across a wide range of applications. Xi et al. [70] discuss various applications involving single agents, multi-agent systems, and human-agent interactions. Wang et al. [71] categorize applications into three key areas: social sciences, natural sciences, and engineering, illustrating the transformative potential of LLM-based agents in empowering these fields. Xie et al. [72] further categorize agents by their application scenarios, highlighting areas such as robotics and embodied AI, where agents make real-time decisions based on multimodal inputs like images, sensor data, and text. Additionally, as noted by Xi et al. [70] and Xie et al. [72], LLM-based agents are becoming increasingly prominent in scientific research, automating tasks such as experiment design, planning, and execution.
@@
To ensure the robustness and reliability of LLM-based agents, future research must focus on overcoming the current limitations. Xie et al. [72] and Shi et al. [68] emphasize the need to develop more robust multimodal perception systems, improve LLM inference efficiency, and establish ethical frameworks to guide agent decision-making. Enhancing memory integration and refining prompt design will also be critical to preventing issues like hallucinations and ensuring agents can effectively transfer knowledge across tasks without degradation.
@@ 4.9 MLLMs in Graph Learning
Multimodal large language models (MLLMs) have been increasingly applied to graph learning tasks, outperforming traditional graph neural networks (GNNs). By incorporating textual attributes and other modalities, MLLMs enhance the representational power of GNNs, enabling improved performance in classification, prediction, and reasoning tasks. Jin et al. proposed a taxonomy categorizing the integration of MLLMs with graphs into enhancers, predictors, and alignment components, highlighting their utility in various graph tasks. Similarly, Chen et al. and Li et al. discuss the integration of LLMs with knowledge graphs, illustrating the benefits of combining graph structures with multimodal data for broader applications [74, 75, 76].
@@ 4.10 Retrieval-Augmented Generation (RAG) in MLLM
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in generating and understanding content across various modalities, such as text, images, and audio. However, their reliance on static training data limits their ability to provide accurate, up-to-date responses, especially in rapidly changing contexts. Retrieval-Augmented Generation (RAG) addresses this issue by dynamically retrieving relevant external information before the generation process. By incorporating real-time and contextually accurate information, RAG enhances both the factuality and robustness of MLLM outputs. In multimodal tasks, RAG not only retrieves textual data but also incorporates multimodal data sources like images and videos, significantly improving the knowledge richness and generation quality of MLLMs [77, 78].
Hu et al. [77] provide a comprehensive survey of RAG’s application in natural language processing, highlighting how retrieving external knowledge during generation effectively mitigates long-tail knowledge gaps and hallucination issues. Zhao et al. [79] further explore how multimodal information retrieval augments generation by improving diversity and robustness. By leveraging information from different modalities, RAG empowers MLLMs to generate more accurate and contextually grounded outputs, especially in tasks requiring cross-modal reasoning such as visual question answering and complex dialogue generation [80, 78].

### v2

@@ 4.3.1 Core Evaluation Domains
This domain evaluates how well MLLMs interpret multimodal inputs, such as text, images, and audio, and integrate information across modalities. Benchmarks in this category include tasks like object detection, scene understanding, and feature extraction. For example, VQA datasets like VQAv2 are foundational for evaluating these abilities but are limited by biases that can inflate model performance [lu2024revisiting].
@@
Higher-level capabilities, such as logical reasoning, problem-solving, and multimodal reasoning, are captured in this domain. Tasks that require more sophisticated reasoning across modalities, like visual question answering with complex scenarios, are included in this category. This domain tests the models’ deeper understanding of the relationships between the modalities [li2024survey].
@@ 4.3.2 Advanced Evaluation Areas
Models need to be evaluated for their robustness, particularly when faced with adversarial prompts or out-of-distribution data. The benchmarks in this category assess how well MLLMs perform under conditions that simulate real-world challenges. Robustness is also critical in ensuring the models’ safety, especially when deployed in domains like autonomous driving or healthcare [chang2024survey, niu2024large]. Furthermore, this area assesses how models manage hallucinations and mitigate biases.
@@
This category includes evaluations focused on specialized domains, such as medical image interpretation or legal text analysis, where models must combine general multimodal abilities with domain-specific knowledge. Benchmarks like TallyQA, which test complex reasoning in specific domains, fall into this category [li2024survey].
@@ 4.3.3 Task-Specific Benchmarks
Early datasets, such as VQAv2, have been widely used to evaluate MLLMs but exhibit limitations due to their susceptibility to language bias and lack of task complexity. In contrast, more recent datasets like TDIUC and DVQA are designed to evaluate models on fine-grained visual understanding, reasoning, and OCR tasks. These datasets provide a more rigorous assessment of the models’ capabilities [lu2024revisiting].
@@
This category focuses on evaluating the interaction between modalities, which is key in tasks like multimodal dialogue or visual question answering that require reasoning across both text and images. Advanced datasets, such as VQDv1, which require complex reasoning and the identification of multiple objects in visual contexts, push models to demonstrate a deeper understanding of multimodal relationships [lu2024revisiting].
@@ 4.3.4 Ethical and Societal Implications
This domain addresses the importance of ensuring that MLLMs do not perpetuate or amplify societal biases. Models must be evaluated on their ability to perform equitably across demographic groups, and benchmarks in this category assess the fairness of the models [li2023surveyfairness].
@@
Ensuring trust in MLLMs is critical, especially in applications where safety is paramount. This category evaluates whether models produce harmful or misleading content and assesses their reliability in sensitive domains. It also evaluates how MLLMs handle uncertainty and avoid hallucinations [gallegos2024bias, bai2024hallucination].
@@ 4.4 Efficiency and Adaptation
The surveys on resource-efficient Multimodal Large Language Models (MLLMs) explore various approaches aimed at reducing computational costs while maintaining performance. Xu et al. [xu2024survey] and Jin et al. [jin2024efficientmultimodallargelanguage] emphasize the growing need for MLLMs to be more accessible, particularly in resource-constrained environments like edge computing. Both surveys provide comprehensive taxonomies, covering advancements in architectures, vision-language integration, training methods, and benchmarks that optimize MLLM efficiency. Key methods discussed include vision token compression, parameter-efficient fine-tuning, and the exploration of transformer-alternative models like Mixture of Experts (MoE) and state space models. These efforts collectively aim to balance computational efficiency with task performance, driving MLLM adoption across various practical applications.
On the other hand, Liu et al. [liu2024fewshotadaptationmultimodalfoundation] address the challenge of adapting MLLMs to specific tasks with limited labeled data. The survey categorizes approaches into prompt-based, adapter-based, and external knowledge-based methods, which help these models generalize better in fine-grained domains such as medical imaging and remote sensing. Few-shot adaptation techniques, such as visual prompt tuning and adapter fine-tuning, are critical for extending the usability of large multimodal models without relying on extensive labeled datasets. Despite advancements, these surveys highlight ongoing challenges, including domain adaptation, model selection, and the integration of external knowledge. Both fields point toward a future where MLLMs are not only more efficient but also more flexible and adaptive in handling diverse, real-world tasks.
@@ 4.5 Data-centric
Bai et al. [bai2024surveymultimodallargelanguage] presents a comprehensive survey on the role of data in the development of Multimodal Large Language Models (MLLMs). The paper emphasizes the importance of the quality, diversity, and volume of multimodal data, which includes text, images, audio, and video, in training MLLMs effectively. It identifies significant challenges in multimodal data collection, such as data sparsity and noise, and explores potential solutions like synthetic data generation and active learning to mitigate these issues. The authors advocate for a more data-centric approach, where the refinement and curation of data take precedence, ultimately improving model performance and advancing MLLM capabilities in an increasingly complex landscape.
Yang et al. [yang2024if] investigates the intersection of large language models (LLMs) and code, framing code as a vital tool that enhances LLMs’ capacity to operate as intelligent agents. The paper discusses how code empowers LLMs in tasks such as automation, code generation, and software development. It also tackles key challenges around code correctness, efficiency, and security, which are essential when deploying LLMs in practical applications. By analyzing the integration of code with LLMs, the authors showcase the potential of these models to evolve into autonomous agents capable of managing complex tasks across various domains, thereby broadening the scope and impact of LLMs in AI-driven innovations.
@@ 4.6 Contunual Learning
Feng et al. [feng2023trends] provides a thorough overview of how Large Language Models (LLMs) are being integrated with external knowledge to enhance their capabilities. The survey categorizes the integration methods into two main approaches: knowledge editing and retrieval augmentation. Knowledge editing involves modifying the input or the model itself to update the outdated or incorrect information, while retrieval augmentation fetches external information during inference without altering the model’s core parameters. The authors present a taxonomy covering these methods, benchmarks for evaluation, and applications such as LangChain and ChatDoctor, which leverage these strategies to address domain-specific challenges. Additionally, the paper explores the handling of knowledge conflicts and suggests future research directions for improving LLM performance in complex, real-world tasks through better integration of multi-source knowledge.
Shi et al. [shi2024continual] offer an extensive survey on the continual learning (CL) of Large Language Models (LLMs), addressing the critical challenges and methodologies in this field. presents an extensive overview of the challenges and techniques related to the continual learning (CL) of Large Language Models (LLMs). The authors focus on two primary directions of continuity: vertical continual learning (adapting models from general to specific domains) and horizontal continual learning (adapting models over time across various domains). They discuss the problems of "catastrophic forgetting," where models lose knowledge of previous tasks, and the complexity of continually updating models to maintain performance on both old and new tasks. The survey outlines key CL methods, including continual pre-training, domain-adaptive pre-training, and continual fine-tuning. It also evaluates various CL techniques, such as replay-based, regularization-based, and architecture-based methods, to mitigate forgetting and ensure knowledge retention. The authors call for more research into evaluation benchmarks and methodologies to counter forgetting and support knowledge transfer in LLMs, making this an underexplored yet crucial area of machine learning research
@@ 4.7 Evaluation Benchmarks
Lu et al. [lu2024revisiting] tackle the key challenges in evaluating multimodal large language models (MLLMs), focusing on the unique complexities these models present. They propose a taxonomy of evaluation metrics designed specifically for multimodal tasks, such as cross-modal retrieval, caption generation, and visual question answering (VQA). The authors highlight how current evaluation frameworks often fail to capture the nuances of multimodal interactions, suggesting the need for more tailored metrics to address these limitations.
Li and Lu [li2024surveybenchmarksmultimodallarge] provide a comprehensive overview of benchmarking datasets and performance metrics in MLLMs. They argue that the absence of standardized protocols across different modalities hinders the fair comparison of models. Their work calls for the establishment of consistent evaluation frameworks that ensure reliable and equitable performance assessment across varied multimodal tasks.
Chang et al. [chang2024survey] examine the evaluation methodologies for large language models (LLMs), emphasizing the importance of moving beyond task-specific performance metrics like knowledge reasoning. They point out that many existing benchmarks, such as HELM and BIG-Bench, overlook critical issues such as hallucinations, fairness, and societal implications. The authors advocate for more holistic evaluation practices that take into account not only the technical capabilities of LLMs but also their trustworthiness, robustness, and ethical alignment in real-world applications. Their work underscores the need to assess both the technical and societal impacts of these models to ensure responsible AI deployment.
@@ 4.8 Agents and Autonomous Systems
Xi et al. [xi2023rise], Wang et al. [wang23surveyonllmautoagents] and a few more papers[xie24lma, gao2023large, xu2024survey] provide comprehensive surveys on the development of LLM-based autonomous agents, highlighting the key modules and frameworks that form the foundation of these systems. At their core, autonomous agents leveraging large language models (LLMs) rely on four essential components: perception, memory, planning, and action. These modules work in synergy to enable agents to perceive their environment, recall previous interactions, and plan and execute actions in real-time. As Xi et al. [xi2023rise] describe, this architecture allows agents to be highly adaptable across a variety of domains, from digital assistants to autonomous vehicles. Wang et al. [wang23surveyonllmautoagents] further emphasize the importance of expanding the perception-action loop by incorporating multimodal inputs, ensuring agents can effectively handle complex real-world scenarios, such as those found in industrial automation and gaming.
@@
Despite these advancements, LLM-based agents face several critical challenges. Both surveys identify knowledge boundaries as a significant issue, where agents are constrained when operating in specialized or underexplored domains. Prompt robustness is another challenge, as even minor changes in prompts can lead to unpredictable or erroneous behavior, including hallucinations, where agents generate false information. Shi et al. [shi2024continual] further explore the challenge of catastrophic forgetting, where agents fail to retain knowledge from previous tasks after being updated with new information. Addressing these challenges requires ongoing improvements in how agents interact with external tools, refine prompt structures, and manage memory to prevent knowledge decay and ensure consistent behavior.
@@
LLM-based agents have demonstrated significant versatility across a wide range of applications. Xi et al. [xi2023rise] discuss various applications involving single agents, multi-agent systems, and human-agent interactions. Wang et al. [wang23surveyonllmautoagents] categorize applications into three key areas: social sciences, natural sciences, and engineering, illustrating the transformative potential of LLM-based agents in empowering these fields. Xie et al. [xie24lma] further categorize agents by their application scenarios, highlighting areas such as robotics and embodied AI, where agents make real-time decisions based on multimodal inputs like images, sensor data, and text. Additionally, as noted by Xi et al. [xi2023rise] and Xie et al. [xie24lma], LLM-based agents are becoming increasingly prominent in scientific research, automating tasks such as experiment design, planning, and execution.
@@
To ensure the robustness and reliability of LLM-based agents, future research must focus on overcoming the current limitations. Xie et al. [xie24lma] and Shi et al. [shi2024continual] emphasize the need to develop more robust multimodal perception systems, improve LLM inference efficiency, and establish ethical frameworks to guide agent decision-making. Enhancing memory integration and refining prompt design will also be critical to preventing issues like hallucinations and ensuring agents can effectively transfer knowledge across tasks without degradation.
@@ 4.9 MLLMs in Graph Learning
Multimodal large language models (MLLMs) have been increasingly applied to graph learning tasks, outperforming traditional graph neural networks (GNNs). By incorporating textual attributes and other modalities, MLLMs enhance the representational power of GNNs, enabling improved performance in classification, prediction, and reasoning tasks. Jin et al. proposed a taxonomy categorizing the integration of MLLMs with graphs into enhancers, predictors, and alignment components, highlighting their utility in various graph tasks. Similarly, Chen et al. and Li et al. discuss the integration of LLMs with knowledge graphs, illustrating the benefits of combining graph structures with multimodal data for broader applications [jin2023large, chen2024knowledge, li2023survey].
@@ 4.10 Retrieval-Augmented Generation (RAG) in MLLM
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in generating and understanding content across various modalities, such as text, images, and audio. However, their reliance on static training data limits their ability to provide accurate, up-to-date responses, especially in rapidly changing contexts. Retrieval-Augmented Generation (RAG) addresses this issue by dynamically retrieving relevant external information before the generation process. By incorporating real-time and contextually accurate information, RAG enhances both the factuality and robustness of MLLM outputs. In multimodal tasks, RAG not only retrieves textual data but also incorporates multimodal data sources like images and videos, significantly improving the knowledge richness and generation quality of MLLMs [hu2024rag, chen2024mllmstrongrerankeradvancing].
Hu et al. [hu2024rag] provide a comprehensive survey of RAG’s application in natural language processing, highlighting how retrieving external knowledge during generation effectively mitigates long-tail knowledge gaps and hallucination issues. Zhao et al. [zhao2023retrievingmultimodalinformationaugmented] further explore how multimodal information retrieval augments generation by improving diversity and robustness. By leveraging information from different modalities, RAG empowers MLLMs to generate more accurate and contextually grounded outputs, especially in tasks requiring cross-modal reasoning such as visual question answering and complex dialogue generation [xu2023large, chen2024mllmstrongrerankeradvancing].